OpenStudy (anonymous):

One-third of eighth graders in Wilson Middle School own a cell phone. If 75 students own cell phones, how many eighth graders are in Wilson Middle School? Choose the correct equation and solution for this situation.

OpenStudy (anonymous):

1/3 + x = 75; x = 225 3x = 75; x = 225 1/3x = 75; x = 225 75/x = 1/3; x= 225

OpenStudy (texaschic101):

1/3x = 75 : x = 225